### v1

# Adiabatic hydrodynamization and quasinormal modes of nonthermal attractors

Abstract
Nonthermal attractors characterize the emergent self-similar evolution of far-from-equilibrium quantum systems, from nuclear collisions to cold-atom experiments. Within the adiabatic hydrodynamization framework, the approach to the nonthermal attractor can be understood as the decay of excited states of an effective Hamiltonian. We use an exactly solvable case of the Boltzmann equation—the longitudinally expanding overoccupied gluon plasma dominated by small-angle elastic collisions—to map the eigenmodes in adiabatic hydrodynamization to the quasinormal mode spectrum of the nonthermal attractor. If generalized, this equivalence may be used as a discovery tool for new phenomena in out-of-equilibrium systems. A byproduct of our analysis is the analytic prescaling solutions for systems undergoing strong longitudinal expansion.
Introduction.– Understanding out-of-equilibrium dynamics and consequent thermalization of quantum many-body systems remains an important subject of contemporary research. Therefore, it is crucial to discover the organizing principles with which to classify and study the emergent dynamics. Prime examples of universal out-of-equilibrium phenomena are nonthermal [1, 2, 3, 4] and hydrodynamic [5, 6, 7] attractors, which are characterized by “memory loss”: a quick collapse of time evolution curves onto a subspace in state space. How a system approaches this attractor behaviour is an active research topic [8, 9, 10, 11, 12, 13, 14, 15, 16]. In particular, some of us have recently employed the Adiabatic Hydrodynamization (AH) [17, 18, 19, 20] and Quasinormal Mode (QNM) [21] frameworks to study the relaxation toward attractor behavior. In this Letter, we establish a correspondence between these two approaches in the context of a longitudinally expanding, overoccupied gluon plasma, relevant for modeling the earliest stages of high-energy nuclear collisions.
Nonthermal attractors or Nonthermal Fixed Points (NTFPs) correspond to attractive regimes of universal self-similar time evolution far-from-equilibrium [1]. They appear in many different branches of physics: in cold atom experiments [22, 23, 24, 25, 26, 27, 28, 29, 30], in the study of the early Universe [31], and in ultra-relativistic heavy-ion collisions [32, 33, 34, 35, 36, 37, 38, 39, 40]. For the latter, at sufficiently high energies, the system is described by a gluon distribution function $f(\tau,p_{z},p_{\perp})$ that depends on proper time ($\tau$), longitudinal ($p_{z}$) and transverse momenta ($p_{\perp}$[39, 40]. The distinction in momentum directions is specific to the heavy-ion case, where a strong longitudinal expansion causes the distribution to be highly anisotropic at early times. In most other cases, the medium is isotropic and all directions are treated equally. A NTFP is defined by the distribution function taking the form
$f(\tau,p_{z},p_{\tau})=\big(\tfrac{\tau}{\tau_{\text{ref}}}\big)^{\alpha}f_{s}\Big(\big(\tfrac{\tau}{\tau_{\text{ref}}}\big)^{\gamma}p_{z},\big(\tfrac{\tau}{\tau_{\text{ref}}}\big)^{\beta}p_{\perp}\Big)\,$, (1)
where $\tau_{\rm ref}$ is some reference time and $f_{s}$ is called the scaling function. When $f$ takes on this form, we say it evolves self-similarly because the functional form of $f_{s}$ remains the same, such that within a long range of subsequent times $f$ can be found by rescaling the amplitude and momenta by powers of time. The scaling function $f_{s}$ together with the scaling exponents $(\alpha,\beta,\gamma)$ characterize the NTFP. This way of characterizing NTFPs also applies to cold atoms experiments [22, 23, 24, 25, 26, 27, 28, 29, 30] and early-Universe cosmology [31], albeit in terms of different variables, with only two scaling exponents $(\alpha,\beta)$ due to isotropy.
We shall demonstrate the correspondence between the AH and QNM frameworks by making use of the Fokker-Planck-type Boltzmann equation [41, 42]
$\frac{\partial f}{\partial\tau}-\frac{p_{z}}{\tau}\frac{\partial f}{\partial p_{z}}=\lambda_{0}I[f]\frac{\partial^{2}f}{\partial p_{z}^{2}}\,$, (2)
with $I=\tfrac{1}{\left(2\pi\right)^{3}}\int d^{3}p\,f^{2}$ and $\lambda_{0}$ a dimensionless parameter characterizing the coupling strength of the theory. This model is exactly solvable down to a single ordinary differential equation, as was first shown by [18] by means of the AH framework. This model reproduces the scaling exponents $(\alpha,\beta,\gamma)=(-2/3,0,1/3)$ predicted by the canonical “bottom-up” thermalization scenario in QCD kinetic theory relevant for heavy-ion collisions in foundational work by Baier, Mueller, Schiff and Son (BMSS) [43, 39, 40], which have since been confirmed by classical-statistical simulations [35, 34]. Furthermore, it matches the shape $f_{s}$ of the distribution function along $p_{z}$, and accounts for the prescaling phenomenon observed in QCD effective kinetic theory (EKT) simulations [10]. Using the AH framework, Ref. [18] demonstrated that it was possible to map the nonlinear Boltzmann equation (2) to a basis in which different components of the distribution function evolve adiabatically, of which the slowest-evolving mode describes the NTFP.
On the other hand, the notion of QNMs has only recently been established in the context of non-expanding, isotropic NTFPs [21]. However, near equilibrium, the notion of QNMs is widespread in usage. In particular, in holographic systems black hole QNMs dictate how a small perturbation to the equilibrium state relaxes back to equilibrium [44, 45, 46] — their analogs in weakly coupled systems are determined by the linearized time evolution operator appearing in the Boltzmann equation. The QNMs of NTFPs provide a different perspective on the emergence and approach to universal out-of-equilibrium dynamics. In this case, they represent self-similar power law corrections to the NTFP solution Eq. (1). If the power laws have negative exponents, the QNMs decay in time, explaining the attractive nature of a NTFP.
Both QNMs and the AH framework can be used to study the approach towards NTFPs, but from different viewpoints. In the AH framework, decay onto the NTFP is encoded in the decay of excited states towards the instantaneous ground state, which asymptotes to the NTFP, while QNMs consider the decay directly to this asymptotic state. If a connection between them may be established across a broad class of systems, one would have a systematic handle to study and solve nonlinear problems in which the information in the linearized response to perturbations may be efficiently used to deduce properties far away from the apparent domain of validity of the linearization. In this Letter, we aim to bridge this gap. Concretely, in the example provided by Eq. (2), we show how the degrees of freedom in the AH picture match those in the QNM picture, and demonstrate their equivalence.
Adiabatic Hydrodynamization.– The AH framework [17, 18, 47, 19, 20] provides an organizing principle to understand the process of memory loss as a system hydrodynamizes starting from an out-of-equilibrium initial condition. In particular, it characterizes the emergence of attractors — both nonthermal fixed points and hydrodynamic attractors. It is based on mapping the Boltzmann equation into a Schrödinger-like equation $\partial_{t}f=-\mathcal{H}(t)f$, where the eigenstates corresponding to eigenvalues with minimal real part are singled out as dynamically preferred configurations. In this way, the emergence of universal dynamics is understood as a consequence of the presence of a gap in the eigenvalue spectrum of $\mathcal{H}$.
In this framework, the organizing principle is adiabaticity: only if the evolution is adiabatic, the low-lying eigenstates are guaranteed to be stable solutions to which the system will be attracted. While adiabaticity is not guaranteed a priori, previous works have shown [18, 19, 20] that it is often possible to obtain an adiabatic description of the dynamics by introducing suitable time-dependent coordinate redefinitions. In particular, Eq. (2) was solved in [18] by introducing arbitrary functions A,B,G{A},{B},{G} that rescale the distribution function and momenta in time: $f(\tau,p_{z},p_{\perp})={A}(\tau){\tilde{f}}(\tau,{\xi},{\chi})$, with ${\xi}\equiv{G}(\tau)p_{z}$ and ${\chi}\equiv{B}(\tau)p_{\perp}$. In terms of the rescaled variables, the kinetic equation (2) takes the form
$\partial_{\tau}{\tilde{f}}$ $=-\frac{{A}^{\prime}(\tau)}{{A}(\tau)}{\tilde{f}}+\Bigg[\frac{1}{\tau}-\frac{{G}^{\prime}(\tau)}{{G}(\tau)}\Bigg]{\xi}\partial_{{\xi}}{\tilde{f}}$
$\quad+\lambda_{0}{A}(\tau)^{2}{G}(\tau){\tilde{I}}\partial^{2}_{{\xi}}{\tilde{f}}$
$=-\mathcal{H}{\tilde{f}}\,$. (3)
For our purposes, we have set $B$ to unity because $p_{\perp}$ does not make an explicit appearance in the equation.
Contrary to quantum mechanics, the effective Hamiltonian $\mathcal{H}$ is time-dependent through its dependence on A{A}, G{G} and ${\tilde{I}}=\tfrac{1}{\left(2\pi\right)^{3}}\int d^{2}{\chi}\int d{\xi}{\tilde{f}}(\tau,{\xi},{\chi})^{2}$. Additionally, it depends on the state on which it acts through $\tilde{I}$}. However, the functions A{A} and G{G} can be chosen such that the eigenstates of the Hamiltonian are time-independent when considered in the rescaled coordinates [18]. If A{A} and G{G} satisfy
$\frac{{G}^{\prime}(\tau)}{{G}(\tau)}$ $=\frac{{A}^{\prime}(\tau)}{{A}(\tau)}+\frac{1}{\tau}\,$, (4)
$\frac{{A}^{\prime}(\tau)}{{A}(\tau)}$ $=-\lambda_{0}{A}(\tau)^{2}{G}(\tau){\tilde{I}}\,$, (5)
then the eigenvalues $\mathcal{E}_{n}$ and left and right eigenstates [^1] A distinction has to be made between left ($\phi^{L}$) and right ($\phi^{R}$) eigenstates because $\mathcal{H}$ is non-Hermitian. $\phi_{n}^{L/R}$ of the Hamiltonian $\mathcal{H}$ are
$\phi_{n}^{L}$ $=\text{He}_{2n}({\xi})\,$, (6)
$\phi_{n}^{R}$ $=\frac{1}{\sqrt{2\pi}(2n)!}\text{He}_{2n}({\xi})e^{-{{\xi}^{2}}/{2}}\,$, (7)
$\mathcal{E}_{n}$ $=2n\frac{{A}^{\prime}(\tau)}{{A}(\tau)}\,$, (8)
and the ensuing evolution is exactly adiabatic. The $\text{He}_{2n}$ denote probabilist’s Hermite polynomials. Concretely, a general solution to Eq. (3) is given by a linear superposition of the eigenstates ${\tilde{f}}={g}({\chi})\sum_{n}a_{n}(\tau)\phi_{n}^{R}({\xi})$, where each eigenstate coefficient evolves independently according to $a_{n}^{\prime}(\tau)=-\mathcal{E}_{n}(\tau)a_{n}(\tau)$, yielding as solutions
$a_{n}(\tau)=a_{n}(\tau_{\rm ref})\left(\frac{{A}(\tau)}{{A}(\tau_{\rm ref})}\right)^{2n}\,.$ (9)
In this way, the non-linear partial differential equation (2) has been reduced to finding A,G{A},{G} that satisfy Eqs. (4) and (5). These in turn enter the eigenvalues and eigenstates above. Since $\tilde{I}$} depends on the initial conditions for all coefficients, a general closed form solution to these equations seems beyond reach. However, this can be achieved in certain cases, which we now examine.
Prescaling in Bjorken flow.– Long before the solutions found in [18], it was realized that Eq. (2) admits a self-similar solution that describes the main features of the BMSS fixed point [35, 39, 40]. However, even in this simpler model, previous works rarely explored how the system relaxes toward this solution starting from arbitrary initial conditions. In QCD EKT simulations [10] it was found that there exists a prescaling regime
$f(\tau,p_{z},p_{\perp})$ $=\bar{A}(\tau)f_{s}(\bar{G}(\tau)p_{z},\bar{B}(\tau)p_{\perp})$
$=\bar{A}(\tau)f_{s}(\bar{\xi},\bar{\chi}),$ (10)
with $\bar{A},\bar{B},\bar{G}$ of a more general form than in Eq. (1). Motivated by this, in [18] it was shown that Eq. (2) exhibits prescaling if the system is initialized on the ground state $\phi_{0}^{R}$ of the spectrum of $\mathcal{H}$. Simultaneously, stability properties of NTFPs were studied in [12]. Recently, [48] showed that prescaling in non-expanding, isotropic NTFPs can be fully captured by a simple time shift, $\tau\to\tau-\tau_{*}$ in Eq. (1), a behavior observed experimentally in a cold-atom system [28]. Prescaling is important because systems may not reach the ideal NTFP form of Eq. (1) within physical timescales [28]. Nevertheless, by allowing generalized rescalings as in Eq. (10), self-similar behavior can still emerge. In heavy-ion collisions, prescaling may play an important role since thermalization occurs on short timescales [39, 40].
The goal of this section is to derive closed form expressions for the functions appearing in the generalized self-similar Ansatz Eq. (10). We replicate the steps carried out in [49] (see Appendix B.2 therein) and illustrate all of the features we will use later on in this Letter — specifically, working toward a QNM analysis.
The functions $\bar{A}$, $\bar{B}$, $\bar{G}$ can be related to each other by considering conservation equations. These can be derived from the Boltzmann equation (2) by taking appropriate moments. In particular, we may consider the transverse energy and number density, given by
$n$ $=\int\frac{d^{3}p}{(2\pi)^{3}}f\,$, (11)
$\epsilon_{\perp}$ $=\int\frac{d^{3}p}{(2\pi)^{3}}p_{\perp}~f\,.$ (12)
Taking these moments of Eq. (2), we immediately find $\partial_{\tau}(\tau n)=0$ and $\partial_{\tau}(\tau\epsilon_{\perp})=0$, because in both cases the weight of the moment is independent of $p_{z}$ and the kernel on the RHS contains derivatives in $p_{z}$, which lead to vanishing boundary terms. Thus,
$\tau n$ $=\tau\bar{A}(\tau)\bar{B}(\tau)^{-2}\bar{G}(\tau)^{-1}n_{s}=\tau_{0}n_{0},$ (13)
$\tau\epsilon_{\perp}$ $=\tau\bar{A}(\tau)\bar{B}(\tau)^{-3}\bar{G}(\tau)^{-1}\epsilon_{\perp,s}=\tau_{0}\epsilon_{\perp,0},$ (14)
with $n_{s}$ and $\epsilon_{\perp,s}$ being the corresponding moments of the scaling function $f_{s}$ in terms of scaled momenta. Here, $\tau_{0}$, $n_{0}$, $n_{s}$, $\epsilon_{\perp,0}$ and $\epsilon_{\perp,s}$ are constants in time. From Eqs. (13) and (14) we see that $\bar{B}(\tau)$ should be a constant $\bar{B}(\tau)=\frac{n_{0}\epsilon_{s}}{n_{s}\epsilon_{0}}\equiv 1$ and $A$ and $G$ are related by
$\bar{G}(\tau)=\frac{\tau}{\tau_{\rm ref}}\bar{A}(\tau).$ (15)
where we defined a reference time $\tau_{\rm ref}$ such that $\tau_{0}n_{0}=\tau_{\rm ref}n_{s}$ and $\tau_{0}\epsilon_{\perp,0}=\tau_{\rm ref}\epsilon_{\perp,s}$. In this way we give the reference time a physical meaning by relating it to the conserved quantities at $\tau=\tau_{0}$ and the ones of the scaling function. Note that Eq. (15) is a solution of Eq. (4) relating $A$ and $G$ in the AH framework.
We can understand Eq. (15) in the following way. The factor $\tau/\tau_{\text{ref}}$ comes from the expansion term in the left-hand side of Eq.(2). The coordinate $(\tau/\tau_{\text{ref}})~p_{z}$ would describe a system that is comoving with expansion. This coordinate transformation is compatible with scaling of the collision kernel we consider. In this comoving frame, $\bar{A}$ and $\bar{G}$ are simply proportional to each other. Here, we can identify a parallel to the isotropic case, where $\bar{A}=\bar{G}^{\sigma}$ [48]. The exponent $\sigma$ is given by the dimensionality $d$ of a system that conserves particle number, while under energy conservation it also depends on the power $z$ of momentum in the dispersion relation of quasiparticles, $\sigma=d+z$. In our case, the Boltzmann equation is effectively a 1D model that describes only dynamics in the longitudinal direction. As such, the conservation laws, Eqs. (13) and (14), are related to the $p_{z}$-profile only. $d=1$ leads to $\bar{A}\propto\bar{G}$ in the frame comoving with expansion, following the logic in the non-expanding, isotropic case.
The self-similar time evolution Ansatz in Eq. (10) implies that $f$ factorizes into a time- and a momentum-dependent part when expressed in terms of the scaled coordinates $\xi$ and $\chi$. Then, the Boltzmann equation (2) for the self-similar Ansatz Eq. (10) can be solved by separation of variables.
$\partial_{\tau}\bar{A}(\tau)=c\bar{A}(\tau)^{4}\tau.$ (16)
$f_{s}+\bar{\xi}\partial_{\bar{\xi}}f_{s}=-\mathcal{\kappa}~\partial^{2}_{\bar{\xi}}f_{s}.$ (17)
Here, $c$ is a separation-of-variables constant and $\kappa=-\lambda_{0}/(\tau_{\text{ref}}~c)\bar{I}_{s}[f_{s}]$, with $\bar{I}_{s}=1/(2\pi)^{3}\int d^{2}\bar{\chi}\int d\bar{\xi}f_{s}^{2}$ the effective temperature factor expressed in terms of scaling quantities. The solutions to these equations are as follows:
$\bar{A}(\tau)=\left(\frac{\tau^{2}\pm\tau_{\ast}^{2}}{\tau_{\rm ref}^{2}}\right)^{-1/3},$ (18)
$f_{s}(\bar{\xi},\bar{\chi})=\bar{g}(\bar{\chi})e^{-\bar{\xi}^{2}/(2\kappa)}.$ (19)
The time shift $\tau_{*}$ in $\bar{A}(\tau)$ emerges naturally as the arbitrary constant of the solution. Also, we chose to define $\tau^{2}_{\rm ref}=-2/(3c)$ to make apparent that for $\tau_{*}=0$ the solution reduces to the BMSS scaling expression $\left(\tau/\tau_{\rm ref}\right)^{\alpha_{\infty}}$ with $\alpha_{\infty}=-2/3$. The equation for $f_{s}$ technically has two solutions, but only one of them integrates to finite values when computing physical observables. Considering $\kappa$ to be a constant, the equation does not have any $\bar{\chi}$-dependence and $g(\bar{\chi})$ is an arbitrary integrable function. However, $\kappa$ depends on $f_{s}$ through $\bar{I}_{s}$. Consistency implies that
$\kappa=\frac{9\lambda_{0}}{2^{8}\pi^{5}}\bar{I}_{\chi}^{2}\tau_{\text{ref}}^{2}$ (20)
with $\bar{I}_{\chi}=\left(\int d^{2}\bar{\chi}~\bar{g}(\bar{\chi})^{2}\right)$. The equation above is important since it shows how $\tau_{\text{ref}}$ enters the scaling function. Also, the above equation means the amplitude $\bar{g}(\bar{\chi})$ of $f_{s}$ is related to $\kappa$, which sets the scale in the Gaussian in Eq. (19). This is in complete agreement with the fact that the rescalings of amplitude and momenta are related, as in Eq. (15).
Eqs. (15), (18) and (19) define the prescaling solution of Eq. (2). In the following we will refer to $\bar{A}(\tau)^{-1}{f}(\tau,\bar{\xi},\bar{\chi})$ as the prescaling frame. In this frame the system becomes effectively static when it is in the prescaling regime. Although beyond our scope, the same solution could be applied to prescaling in the full EKT [49], though the scaling-breaking terms add complexity [48].
QNMs in Bjorken flow.– In [21] it was shown that, near a non-expanding, isotropic NTFP, linear order deviations from it can be written as a superposition of quasinormal modes (QNMs). These decay in a power-law manner to the NTFP with different exponents $\Omega$, which define the QNM frequencies. We generalize this analysis to the expanding system by considering a perturbation in the prescaling frame $\bar{A}(\tau)^{-1}f(\tau,p_{z},p_{\perp})=f_{s}(\bar{\xi},\bar{\chi})+\delta\bar{f}(\tau,\bar{\xi},\bar{\chi})$ with $\delta\bar{f}\ll 1$. At linear order in $\delta\bar{f}$, Eq. (2) yields an equation of the form
$\frac{\bar{A}(\tau)}{\bar{A}^{\prime}(\tau)}{\partial_{\tau}}\delta\bar{f}(\tau,\bar{\xi},\bar{\chi})=\hat{O}[f_{s}](\bar{\xi})\delta\bar{f}(\tau,\bar{\xi},\bar{\chi}).$ (21)
The solutions are the QNMs of the NTFP:
$\delta\bar{f}(\tau,\bar{\xi},\bar{\chi})=\delta\cdot\bar{A}(\tau)^{\Omega}f_{\Omega}(\bar{\xi},\bar{\chi}),$ (22)
where $\delta\ll 1$. Note that in [21] the QNMs were defined with $i\Omega$ to be in line with the convention that a negative imaginary part indicates decay. We choose to omit this here to simplify notation.
Similarly to the non-expanding case, the QNM amplitude is a power of the time-dependent amplitude of the prescaling expression, Eq.(10). At late times, the QNMs approach the NTFP as power laws in time; see Eq. (18). This similarity is not surprising if we realize that we consider the system in rescaled coordinates such that it looks effectively static. In other words, the expansion has been taking care of on the level of the rescaling of amplitude and momenta.
The momentum dependence is determined by an eigenvalue equation,
$\Omega f_{\Omega}=\hat{O}[f_{s}]f_{\Omega}\\ =-f_{\Omega}-\bar{\xi}\partial_{\bar{\xi}}f_{\Omega}-\kappa[f_{s}]\partial^{2}_{\bar{\xi}}f_{\Omega}-\kappa\frac{\bar{I}_{\delta}[f_{s},f_{\Omega}]}{\bar{I}_{s}[f_{s}]}\partial_{\bar{\xi}}^{2}f_{s},$ (23)
where $\bar{I}_{\delta}=1/(2\pi)^{3}\int d^{2}\bar{\chi}\int d\bar{\xi}2f_{s}f_{\Omega}$. For simplicity, we will in the following assume all perturbations to have the same transverse momentum distribution as the background, which we alleviate in App. A. Eq. (23) can be solved analytically, resulting in the following QNM frequencies and modes.
$\Omega_{1}=3,\ \Omega_{n}=2n\ \text{for}\ n=0,2,3,4,$... (24)
$\begin{split}&f_{\Omega_{1}}(\bar{\xi},\bar{\chi})\propto\bar{g}(\bar{\chi})e^{-{\bar{\xi}^{2}}/({2\kappa})}\text{He}_{2}(\bar{\xi}/\sqrt{\kappa})\\ &f_{\Omega_{n}}(\bar{\xi},\bar{\chi})\propto\bar{g}(\bar{\chi})e^{-{\bar{\xi}^{2}}/({2\kappa})}\Bigg[\text{He}_{2n}(\bar{\xi}/\sqrt{\kappa})\\ &\qquad\qquad+\frac{\bar{I}_{\text{He}_{2n}}}{\sqrt{\pi}\sqrt{\kappa}(\Omega_{n}-3)}\text{He}_{2}(\bar{\xi}/\sqrt{\kappa})\Bigg],\ n\neq 1.\end{split}$ (25)
Here, $\bar{I}_{\text{He}_{2n}}=\int d\bar{\xi}e^{-{\bar{\xi}^{2}}/\kappa}\text{He}_{2n}(\bar{\xi}/\sqrt{\kappa})$ is the longitudinal momentum factor of the integral $\bar{I}$ for the perturbation.
Physical interpretation of QNMs.– The first two modes can be understood in physical terms by making connection with the prescaling notion. As we will now elaborate, shifting $\tau_{\rm ref}$ and $\tau_{*}$ lead to the zero mode $\Omega_{0}=0$ and to $\Omega_{1}=3$, respectively. We will refer to $\Omega_{1}=3$ as the prescaling QNM in the following, since a non-zero $\tau_{*}$ leads to prescaling. We will showcase how the prescaling mode arises under a shift $\tau_{*}\rightarrow\tau_{*}+\delta\tau_{*}$. The effects on the prescaling functions $\bar{A}$ and $\bar{\xi}=\bar{G}(t)p_{z}$ is as follows [see Eqs. (18) and (15)]:
$\begin{split}&\bar{A}(\tau)\xrightarrow{\tau_{*}\rightarrow\tau_{*}+\delta\tau_{*}}\bar{A}(\tau)\pm\alpha_{\infty}\frac{\tau_{*}\delta\tau_{*}}{\tau_{\rm ref}^{2}}\bar{A}(\tau)^{4}\,,\\ &\bar{\xi}\xrightarrow{\tau_{*}\rightarrow\tau_{*}+\delta\tau_{*}}\bar{\xi}\pm\alpha_{\infty}\frac{\tau_{*}\delta\tau_{*}}{\tau_{\rm ref}^{2}}\bar{A}(\tau)^{3}\bar{\xi}.\end{split}$ (26)
Under these changes $\bar{A}(\tau)f_{s}(\bar{\xi})$ changes as
$\bar{A}(\tau)f_{s}(\bar{\xi})\xrightarrow{\tau_{*}\rightarrow\tau_{*}+\delta\tau_{*}}\bar{A}(\tau)f_{s}(\bar{\xi})\\ \pm\alpha_{\infty}\frac{\tau_{*}\delta\tau_{*}}{\tau_{\rm ref}^{2}}\bar{A}(\tau)^{4}\left(f_{s}(\bar{\xi})+\bar{\xi}\partial_{\bar{\xi}}f_{s}(\bar{\xi})\right).$ (27)
For our solution of the scaling function Eq. (19), this indeed results in the QNM solution with $\Omega_{1}=3$, see Eq. (25).
The analogous analysis for $\tau_{\rm ref}\rightarrow\tau_{\rm ref}+\delta\tau_{\rm ref}$ leads to the QNM solution with $\Omega_{0}=0$. Note that in this case one needs to take into account the explicit dependence of $f_{s}$ on $\tau_{\rm ref}$. This enters through $\kappa\propto\tau_{\text{ref}}^{2}$, see Eq. (20).
Finally, we show that in order for the prescaling frame to extend to QNMs, $\delta n=0$ and $\delta\epsilon_{\perp}=0$ need to hold. For the perturbed system conservation of particle number leads to:
[…]
Comparing to the unperturbed case, Eq.(13), gives $\delta n(\tau)=0$. The reasoning for transverse energy density $\epsilon_{\perp}$ is the same. All the QNMs above, except for $\Omega_{0}=0$, satisfy this constraint. We thus identify the zero mode with perturbations in our conserved quantities. This is consistent with the zero mode arising from a shift in $\tau_{\rm ref}$, since the latter is related to the conserved combination $\tau_{0}n_{0}$ and $\tau_{0}\epsilon_{\perp 0}$.
Importantly, the above means the prescaling frame, defined by the expressions Eq. (15) and (18), is solely defined on the level of the scaling function $f_{s}$. We can thus use the same $\bar{A}$ and $\bar{G}$ when considering the QNMs on top of $f_{s}$.
Nonlinear solution in terms of QNMs.– The QNMs are by definition solutions to the linearized Boltzmann equation in the overoccupied limit, Eq. (21). In this section we show how, starting from the QNM solutions, we can construct solutions to the non-linear equation, Eq. (2), that are accurate to arbitrary order in the perturbation. We start by considering Eq. (2) in the prescaling frame: $f(\tau,p_{z},p_{\perp})=\bar{A}(\tau)\bar{f}(\tau,\bar{\xi},\bar{\chi})$.
$\frac{\bar{A}(\tau)}{\bar{A}^{\prime}(\tau)}\partial_{\tau}\bar{f}+\bar{f}+\bar{\xi}\partial_{\bar{\xi}}\bar{f}+\kappa\frac{\bar{I}(\tau)}{\bar{I}_{s}}\partial^{2}_{\bar{\xi}}\bar{f}=0$ (29)
where we explicitly denote the time dependence of $\bar{I}(\tau)=\frac{1}{\left(2\pi\right)^{3}}\int d^{2}\bar{\chi}\int d\bar{\xi}\bar{f}^{2}$ in contrast to the scaling version $\bar{I}_{s}$. Thus far we have just reparameterized our problem, so nothing has changed physically.
To construct a solution of this equation, we start from a linear combination of QNMs on top of the scaling function, which solves it to linear order: $\bar{f}_{0}=f_{s}+\sum_{n}c_{n}\bar{A}(\tau)^{\Omega_{n}}f_{\Omega_{n}}$, where $|c_{n}|\bar{A}(\tau)^{\Omega_{n}}\ll 1$. A solution to any given order in the perturbation can be found by expanding in the coefficients $c_{n}\bar{A}(\tau)^{\Omega_{n}}$. Eq. (29) will dictate the appropriate momentum dependence of higher order terms. In Appendix B, we explain this procedure in more detail.
Here, we give an example. For simplicity we choose $\kappa=1$ by fixing the amplitude of $\bar{g}(\bar{\chi})$, see also Eq. (20). Consider only the two lowest QNMs, then the solution to Eq. (29) to second order in the $c_{n}$ is
$\begin{split}\bar{f}&=\bar{g}(\bar{\chi})\Bigg[e^{-\bar{\xi}^{2}/2}+c_{1}\bar{A}(\tau)^{3}e^{-\bar{\xi}^{2}/2}\left(-1+\bar{\xi}^{2}\right)\\ &+c_{2}\bar{A}(\tau)^{4}e^{-\bar{\xi}^{2}/2}\left(18-30\bar{\xi}^{2}+4\bar{\xi}^{4}\right)\\ &+\frac{1}{2}c_{1}^{2}\bar{A}(\tau)^{6}e^{-\bar{\xi}^{2}/2}\left(4-7\bar{\xi}^{2}+\bar{\xi}^{4}\right)\\ &+2c_{1}c_{2}\bar{A}(\tau)^{7}e^{-\bar{\xi}^{2}/2}\left(-45+114\bar{\xi}^{2}-33\bar{\xi}^{4}+2\bar{\xi}^{6}\right)\\ &+\frac{3}{5}c_{2}^{2}\bar{A}(\tau)^{8}e^{-\bar{\xi}^{2}/2}\left(803-2093\bar{\xi}^{2}+630\bar{\xi}^{4}-40\bar{\xi}^{6}\right)\Bigg].\end{split}$ (30)
Similar nonlinear solutions have been considered in the context of black hole QNMs [50, 51]. There transient effects are studied, based on the fact that the QNMs are the eigenvalues of a non-normal operator. We anticipate that aspects of these findings could apply in the context of NTFP QNMs. To this end, a more careful investigation of the non-normality of the linearized operator in Eq. (23) is necessary.
Comparison to Adiabatic Hydrodynamization results.– Having constructed the QNM solutions in the linear and nonlinear regimes, we are now in a position to make an explicit connection with the results from the AH framework. While QNMs describe the decay directly onto the NTFP, the excited states of AH quantify the decay to the ground state, which itself evolves towards the NTFP via the evolution of ${A}(\tau)$. To connect these pictures, we need to examine the behavior of ${A}(\tau)$ close to the scaling solution as given by Eqs. (4) and (5). No general solution exists but we can construct one perturbatively in the excited states. We show that the resulting solution is exactly the non-linear solution in terms of QNMs we found above, Eq. (30).
To showcase this, consider a small perturbation $\propto\phi_{n}^{R}$ on top of the right-handed ground state $\phi_{0}^{R}$: ${\tilde{f}}=\phi_{0}^{R}({\xi})+\delta_{n}\bar{A}(\tau)^{2n}\phi^{R}_{n}({\xi})$, with $\delta_{n}\ll 1$. For $\delta_{n}=0$, we can match the solution to our prescaling solution, Eqs. (18) and (19) with $\kappa=1$. To this end, we relate A{A} and G{G} as in Eq. (15) and choose the amplitudes in $\bar{g}(\bar{\chi})$ and ${g}({\chi})$ such that
$\begin{split}&\bar{I}_{s}=\frac{2}{3\lambda_{0}\tau_{\rm ref}}={\tilde{I}}[{\tilde{f}}=\phi_{0}^{R}].\end{split}$ (31)
Note that this means $\bar{g}(\bar{\chi})=\sqrt{2\pi}~{g}({\chi})$.
Now, by means of Eqs. (4) and (5), we are able to calculate A{A} as a series expansion in $\delta_{n}$: ${A}=\bar{A}+\sum_{i}^{\infty}(\delta_{n})^{i}~\bar{A}_{i}$. We will consider here the case $n=2$. Plugging the resulting solution into ${\tilde{f}}=\phi_{0}^{R}({\xi})+\delta_{2}\bar{A}(\tau)^{2}\phi^{R}_{2}({\xi})$ and truncating after $O(\delta_{2}^{2})$, we find
$\begin{split}{\tilde{f}}&=\bar{g}(\bar{\chi})\Bigg[e^{-\bar{\xi}^{2}/2}-(d_{1}+d_{2}\delta_{2})\delta_{2}\bar{A}(\tau)^{3}e^{-\bar{\xi}^{2}/2}\left(-1+\bar{\xi}^{2}\right)\\ &+\frac{1}{24}\delta_{2}\bar{A}(\tau)^{4}e^{-\bar{\xi}^{2}/2}\left(18-30\bar{\xi}^{2}+4\bar{\xi}^{4}\right)\\ &+\frac{1}{2}d_{1}^{2}\delta_{2}^{2}\bar{A}(\tau)^{6}e^{-\bar{\xi}^{2}/2}\left(4-7\bar{\xi}^{2}+\bar{\xi}^{4}\right)\\ &-\frac{1}{12}d_{1}\delta_{2}^{2}\bar{A}(\tau)^{7}e^{-\bar{\xi}^{2}/2}\left(-45+114\bar{\xi}^{2}-33\bar{\xi}^{4}+2\bar{\xi}^{6}\right)\\ &+\frac{1}{960}\delta_{2}^{2}\bar{A}(\tau)^{8}e^{-\bar{\xi}^{2}/2}\left(803-2093\bar{\xi}^{2}+630\bar{\xi}^{4}-40\bar{\xi}^{6}\right)\Bigg],\end{split}$ (32)
where we made use of ${\xi}={G}(\tau)p_{z}$ and Eq. (4). $d_{1}$ and $d_{2}$ are integration constants that are picked up order by order in the solution for A{A}.
Comparing this to Eq. (30), we find that there is an exact match when considering terms to $O(\delta_{2}^{2})$ after identifying $c_{1}=-(d_{1}+d_{2}\delta_{2})\delta_{2}$ and $c_{2}=\delta_{2}/24$. This serves as a non-trivial cross-check of the earlier solution in Eq. (30), since this approach is agnostic to the expressions for the QNMs. More importantly, it is now clear how the QNMs enter the AH solution. The first excited state excites the second QNM $\Omega_{2}=4$ and the prescaling QNM $\Omega_{1}=3$. Notably, no contributions from other QNMs enter the solution. This structure persists when considering additional excited states on top of the ground state. The effect of the $n$th excited state can be captured by the prescaling and the (n1)(n-1)th QNMs.
Therefore, we achieved our goal of understanding the explicit time dependence of A{A} close to the NTFP. Indeed the approach is captured by a series of terms that decay according to different power-laws. This is the QNM picture we presented before. Importantly, the slowest power law is $\bar{A}^{3}$ instead of $\bar{A}^{2}$, which one might have naively expected from the AH picture. From the AH perspective, this is due to the fact that perturbations in the evolution of $A$ in terms of $\delta_{n}$, in addition to the perturbations themselves, generate a change to the coordinate frame on which they propagate, leading to a power law that is not directly determined by the eigenvalues $\mathcal{E}_{n}$. Most importantly, we have established the equivalence between the two pictures and how to construct the mapping.
Outlook.– In this Letter, we have found explicit solutions for universal features of the dynamics of nonlinear systems, as well as organizing principles with which to generalize these findings. These give us a platform to develop concrete, rigorous way of understanding out-of-equilibrium phenomena. Specifically, by connecting the explicit QNM solutions around a NTFP to the exact (albeit implicit) nonlinear solutions of the dynamics of the same theory, we have opened the gates to a systematic exploration of the nonlinear response regime around NTFPs and the general relaxation dynamics towards them.
Our results explicitly show how to extend QNMs into the nonlinear regime. Furthermore, we have shown how these solutions can be matched to fully nonlinear solutions obtained in the AH framework. In tandem, these two developments show how one may seek and find exact adiabatic solutions that extend non-perturbatively away from the NTFP by using the information encoded in the QNMs. Here, we have focused on the BMSS fixed point, but we speculate that the steps we took to connect these two concepts may be applicable for a broad class of NTFPs. Establishing the breadth of this class — which our work exemplifies — is an exciting new open problem, with potential to give us deep insights about the general classification and structure of universal out-of-equilibrium phenomena.
In this work, we have considered a simple idealized model, where a large anisotropy justifies the complete decoupling of the transverse sector. A natural next step would be to study how our results extend to models that go beyond our simplifying assumptions by including transverse dynamics or even fully dropping the assumption of longitudinal momentum domination. As well as for models that have a dual cascade [52, 53], making the connection stronger to experimental NTFPs that observe such dual cascades [28, 26, 24, 25, 27, 29, 30]. On an even broader scope, as the concept of QNMs originates from holography, recent insights about their role in that context [54, 55, 56, 57, 51, 50] might synergize with the results of this paper to gain an even deeper understanding of the process. Similarly, important insights might be gained from exploring a connection to singulants [58], which were introduced as an extension of the notion of QNMs beyond near-equilibrium systems, motivated by the large order behavior of the factorially divergent hydrodynamic gradient expansion.
Acknowledgements.– We would like to thank Thimo Preis for his contributions to the study of prescaling in expanding nonthermal attractors during the early stages of this project. We thank Yi Yin for useful comments on the draft. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant number: 101089093 / project acronym: High-TheQ). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This work was partially supported by the Priority Research Area Digiworld under the program Excellence Initiative - Research University at the Jagiellonian University in Krakow. The work of BSH was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP) and by grant 994312 from the Simons Foundation. The work of AM is supported by the DFG through Emmy Noether Programme (project number 496831614) and through CRC 1225 ISOQUANT (project number 27381115). We thank ECT* for support at the workshop “Attractors and thermalization in nuclear collisions and cold quantum gases,” where we profited from helpful discussions.
Appendix A QNM solution with additional transverse momentum dependence
Here we give the solutions to Eq. (23) considering perturbations with transverse momentum dependence that is different than that of the scaling function $\bar{g}(\bar{\chi})$:
$\Omega_{1}=3,\ \Omega_{n}=2n\ \text{for}\ n=0,2,3,4,$... (33)
$\begin{split}&f_{\Omega_{1}}(\bar{\xi},\bar{\chi})\propto\bar{g}(\bar{\chi})e^{-{\bar{\xi}^{2}}/({2\kappa})}\text{He}_{2}(\bar{\xi}/\sqrt{\kappa})\\ &f_{\Omega_{n}}(\bar{\xi},\bar{\chi})\propto e^{-{\bar{\xi}^{2}}/({2\kappa})}\Bigg[\bar{g}_{n}(\bar{\chi})\text{He}_{2n}(\bar{\xi}/\sqrt{\kappa})+\bar{g}(\bar{\chi})\frac{\bar{I}_{\text{He}_{2n}}\bar{I}_{\chi n}}{\sqrt{\pi}\sqrt{\kappa}\bar{I}_{s\chi}(\Omega_{n}-3)}\text{He}_{2}(\bar{\xi}/\sqrt{\kappa})\Bigg],\ n\neq 1.\end{split}$ (34)
The additional transverse momentum dependence $\bar{g}_{n}(\bar{\chi})$ is an arbitrary function so long as it describes a physical (i.e., positive) distribution function, $\bar{I}_{s\chi}=\int d^{2}\bar{\chi}\bar{g}(\bar{\chi})^{2}$ is the transverse momentum factor of the integral $\bar{I}_{s}$, and $\bar{I}_{\chi n}=\int d^{2}\bar{\chi}\bar{g}(\bar{\chi})\bar{g}_{\delta,n}(\bar{\chi})$ is the same integral for the perturbation, with the corresponding longitudinal momentum factor $\bar{I}_{\text{He}_{2n}}=\int d\bar{\xi}e^{-{\bar{\xi}^{2}}/\kappa}\text{He}_{2n}(\bar{\xi}/\sqrt{\kappa})$.
Appendix B Scaling perturbation to arbitrary order
Starting from a linear combination of the zeroth and first order result,
$\bar{f_{0}}=f_{s}+\sum_{n}c_{n}\bar{A}(\tau)^{\Omega_{n}}f_{\Omega_{n}},$ (35)
we want to construct solutions to the equation
$\frac{\bar{A}(\tau)}{\bar{A}^{\prime}(\tau)}\partial_{\tau}\bar{f}+\bar{f}+\bar{\xi}\partial_{\bar{\xi}}\bar{f}+\mathcal{\kappa}\frac{\bar{I}}{\bar{I}_{s}}\partial^{2}_{\bar{\xi}}\bar{f}=0$ (36)
to arbitrary order in the coefficients $c_{n}$.
When plugging Eq. (35) into Eq. (36), since QNMs solve the equation up to linear order, only terms that are quadratic and cubic in the QNMs remain. These can be countered by higher order terms. By iteration, the equation can be solved order by order in an expansion of products of the different $c_{n}\bar{A}(\tau)^{\Omega_{n}}$. In the end, one only has to solve for the remaining momentum dependence. For example, the term $\propto c_{1}^{2}$ is proportional to $\bar{A}(\tau)^{2\Omega_{1}}=\bar{A}(\tau)^{6}$. By adding a term $c_{1}^{2}\bar{A}(\tau)^{2\Omega_{1}}P(\bar{\xi})f_{s}$ to $\bar{f}_{0}$ we can solve for $P(\bar{\xi})$. $P(\bar{\xi})$ will be a polynomial, as shown below. These steps can be iterated to find higher order corrections. We now lay out how to obtain the solution to arbitrary order in the coefficients $c_{n}$.
Following the arguments above, we can construct, starting from $f_{0}$, a solution to the evolution equation via the Ansatz
$\bar{f}$ $=\bar{f}_{0}+\sum_{|\alpha|\geq 2}\mathbf{c}^{\alpha}\bar{A}(\tau)^{\mathbf{\Omega}\cdot\alpha}P_{\alpha}(\bar{\xi})f_{s}$ (37)
$=\sum_{|\alpha|\geq 0}\mathbf{c}^{\alpha}\bar{A}(\tau)^{\mathbf{\Omega}\cdot\alpha}P_{\alpha}(\bar{\xi})f_{s}$ (38)
using multi-index notation for $\alpha$. The second equality defines $P_{0}=1$ and the $P_{\alpha}$ with $\alpha_{i}=\delta_{in}$ as the corresponding combinations of Hermite polynomials in the $n$-th QNM, respectively. Plugging this into Eq. (36), coefficient comparison in monomials of the $c_{n}$ yields a set of equations that relate the $P_{\alpha}(\bar{\xi})$ and allow to compute them order by order. If all $P_{\beta}$ for $\beta\leq\alpha$, $\beta\neq\alpha$ are known, $P_{\alpha}$ is obtained as the solution of the following integro-differential equation (setting $\kappa=1$):
$(\mathbf{\Omega}\cdot\alpha)P_{\alpha}(\bar{\xi})-\bar{\xi}P_{\alpha}^{\prime}(\bar{\xi})+P_{\alpha}^{\prime\prime}(\bar{\xi})+2(\bar{\xi}^{2}-1)\int\frac{d{\bar{\xi}^{\prime}}}{\sqrt{\pi}}P_{\alpha}({\bar{\xi}^{\prime}})f_{s}^{2}({\bar{\xi}^{\prime}})$
$=\sum_{\begin{subarray}{c}0\neq\beta+\gamma\leq\alpha\\ \beta,\gamma\neq\alpha\end{subarray}}[(1-\bar{\xi}^{2})P_{\gamma}(\bar{\xi})+2\bar{\xi}P_{\gamma}^{\prime}(\bar{\xi})-P_{\gamma}^{\prime\prime}(\bar{\xi})]\int\frac{d{\bar{\xi}^{\prime}}}{\sqrt{\pi}}P_{\alpha-\beta-\gamma}({\bar{\xi}}^{\prime})P_{\beta}({\bar{\xi}^{\prime}})f_{s}^{2}({\bar{\xi}^{\prime}}).$ (39)
Assuming further that the $P_{\alpha}$ are of the form $P_{\alpha}(\bar{\xi})=\sum_{i}c_{\alpha,i}\bar{\xi}^{2i}$, a second coefficient comparison yields algebraic equations that allow to compute the $c_{\alpha,i}$:
$(\mathbf{\Omega}\cdot\alpha-2i)c_{\alpha,i}+(2i+2)(2i+1)c_{\alpha,i+1}+2(\delta_{i1}-\delta_{i0})\sum_{j}c_{\alpha,j}\pi^{-1/2}{\Gamma(j+1/2)}$
$=\sum_{\begin{subarray}{c}0\neq\beta+\gamma\leq\alpha\\ \beta,\gamma\neq\alpha\end{subarray}}[(4i+1)c_{\gamma,i}-c_{\gamma,i-1}-(2i+2)(2i+1)c_{\gamma,i+1}]\sum_{j,k}c_{\alpha-\beta-\gamma}c_{\beta}\pi^{-1/2}\Gamma(j+k+1/2).$ (40)